OpenStudy (gagecool):

What is the y-intercept of the graph of the line given by the equation 3y + 4x = 15? A. 5 B. 4 C. 3 D. -4/3

OpenStudy (qwertty123):

Okay do you know what a y intercept is?

OpenStudy (priyar):

rewrite this eq in the form of: x/a+y/b=1

OpenStudy (gagecool):

yes

OpenStudy (priyar):

where a and b are the x and y intercepts resp.

OpenStudy (qwertty123):

Which number is on the y axes?

OpenStudy (gagecool):

3

OpenStudy (qwertty123):

Okay. Then that is your answer. :)

OpenStudy (gagecool):

really i would have never guessed that thanks gave you a medal

OpenStudy (priyar):

no thats not the answer

OpenStudy (qwertty123):

Then explain it?

OpenStudy (gagecool):

wait that not the right answer then i am confused

OpenStudy (priyar):

first write the eq in the form i told u before

OpenStudy (gagecool):

i did that but i dont see how that helps so please explain

OpenStudy (priyar):

what eq did u get can u write it here?

OpenStudy (qwertty123):

Sorry I am wrong. The wrong place to be sorry! I know this answer! I am so sorry @Gagecool I answered for the wrong person!

OpenStudy (qwertty123):

In this formula : y tells us how far up the line goes x tells us how far along m is the Slope or Gradient i.e. how steep the line is b is the Y-intercept i.e. where the line crosses the Y axis The X and Y intercepts and the Slope are called the line properties. We shall now graph the line 3y-4x-15 = 0 and calculate its properties Graph of a Straight Line : Calculate the Y-Intercept : Notice that when x = 0 the value of y is 5/1 so this line "cuts" the y axis at y= 5.00000 y-intercept = 15/3 = 5 Calculate the X-Intercept : When y = 0 the value of x is 15/-4 Our line therefore "cuts" the x axis at x=-3.75000 x-intercept = 15/-4 = -3.75000 Calculate the Slope : Slope is defined as the change in y divided by the change in x. We note that for x=0, the value of y is 5.000 and for x=2.000, the value of y is 7.667. So, for a change of 2.000 in x (The change in x is sometimes referred to as "RUN") we get a change of 7.667 - 5.000 = 2.667 in y. (The change in y is sometimes referred to as "RISE" and the Slope is m = RISE / RUN) Slope = 2.667/2.000 = 1.333

OpenStudy (gagecool):

oh thats okay

OpenStudy (priyar):

@Gagecool did u write it in that form?

OpenStudy (priyar):

first div by 15 on both sides

OpenStudy (qwertty123):

Here this goes to graph the strait line above!

OpenStudy (gagecool):

4x + 3y = 15 that what i got by writing it like that i think that is right

OpenStudy (qwertty123):

Notice that when x = 0 the value of y is 5/1 so this line "cuts" the y axis at y= 5.00000 y-intercept = 15/3 = 5 Calculate the X-Intercept : When y = 0 the value of x is 15/-4 Our line therefore "cuts" the x axis at x=-3.75000 x-intercept = 15/-4 = -3.75000 Calculate the Slope : Slope is defined as the change in y divided by the change in x. We note that for x=0, the value of y is 5.000 and for x=2.000, the value of y is 7.667. So, for a change of 2.000 in x (The change in x is sometimes referred to as "RUN") we get a change of 7.667 - 5.000 = 2.667 in y. (The change in y is sometimes referred to as "RISE" and the Slope is m = RISE / RUN) Slope = 2.667/2.000 = 1.333 Geometric figure: Straight Line Slope = 2.667/2.000 = 1.333 x-intercept = 15/-4 = -3.75000 y-intercept = 15/3 = ?

OpenStudy (qwertty123):

Y-intercept equals?

OpenStudy (priyar):

no @Gagecool thats not it...can we go one step at a time?

OpenStudy (qwertty123):

Sorry to confuse you @Gagecool ! Hope this helps.

OpenStudy (gagecool):

When i divide 15 on both sides i get 5 on one and 3.75 on the other what do i do next

OpenStudy (qwertty123):

correct!

OpenStudy (gagecool):

that okay Qwertty123

OpenStudy (qwertty123):

@Gagecool thanks for understanding! :D

OpenStudy (priyar):

ya.. u get: \[3y/15 + 4x/15=1\]

OpenStudy (priyar):

Now bring the coeffecient below 15 for both x and y

OpenStudy (gagecool):

the answer comes out to be 5 when i divide then add

OpenStudy (priyar):

yes

OpenStudy (qwertty123):

Yes you are correct! Way to go @Gagecool

OpenStudy (gagecool):

so the answer is 5

OpenStudy (qwertty123):

Yes you did it!!!

OpenStudy (priyar):

\[y/(15/3) + x/(15/4) =1\]

OpenStudy (priyar):

the denominator below y which is 5 (on reducing) is the y intercept

OpenStudy (gagecool):

then thank you both for helping me out and giving me the answer 5.

OpenStudy (priyar):

the denominator below x is the x-intercept...Thats y i asked u to write in this form!

OpenStudy (gagecool):

oh okay i see

OpenStudy (qwertty123):

You welcome @Gagecool !!!

OpenStudy (priyar):

if u r given such q's first write it in the above form...it will be very helpful!

OpenStudy (priyar):

welcome @Gagecool

OpenStudy (qwertty123):

@priyar you did a great job helping! I think you also deserve a medal! :D I will give you one! :)

OpenStudy (priyar):

thank u!

OpenStudy (gagecool):

yeah i was able to give you both one

OpenStudy (qwertty123):

@Gagecool I will become your fan since I already gave my medal away.

OpenStudy (gagecool):

well thank you, you think you can help me out on this next question

OpenStudy (qwertty123):

Um not sure... let me see the question :)

OpenStudy (gagecool):

okay i will put it in the open question box above so get ready to see it

OpenStudy (qwertty123):

okay :D